# zkSTAR: A zero knowledge system for time series attack detection enforcing regulatory compliance in critical infrastructure networks

### v1

@@ IV-B Temporal Consistency
The state space system is considered to be temporally consistent if for each consecutive time step pairs $t,t+1$, the following condition holds.
@@
As a consequence of Definition 7 we can make the temporal consistency claim presented in Theorem IV.1.
Theorem IV.1 (zk-STAR Temporal Consistency Conditions).
@@
Proof provided in Appendix -C
Theorem IV.1 shows that the integrity of temporal posteriori predictions from an EKF driven state space system can be established using a zkSNARK proof artifact if and only if the verification keys remain invariant; the hash of the inputs to the EKF model at each time step is equal to the hash of the previous EKF output; and the proof artifact itself can be verified using the verification keys. In other words, Theorem IV.1 demonstrates that if temporally inconsistent values for carrying out the EKF predictions, then either the verification fails for at least one time step or the keys used for verifying the proof artifact will differ. Therefore, to establish temporally consistency from a regulatory perspective, it is necessary and sufficient to check for the invariance of keys, alignment of inputs and outputs and the verification status across successive time steps.
IV-C Statistical Consistency
Consider the local attack detection framework characterized by the $\chi^{2}$ hypothesis test presented in Section III-B with a Type-I and Type-II error rates of $\alpha$ and $\beta$ respectively. The $\chi^{2}$ test relies on the residual and state space estimates provided by $\mathcal{M}$ that is parametrized by $\theta$. As a result of Equation (34), the local attack detection framework is expected to yield an alarm value denoted by $\rho_{t}=1$, when an attack is taking place while maintaining $\rho_{t}=0$ under normal circumstances. We utilize these fundamental principles to establish the notion of statistical consistency. For the purposes of statistical consistency, we denote the genuine and the reported outcomes of the $\chi^{2}$ test as $\rho^{real}_{t}\in\{0,1\}$ and $\rho^{rep}_{t}\in\{0,1\}$ respectively.
@@ IV-C Statistical Consistency
Statistical consistency ensures that reported alarm values are consistent with the outcomes of the attack detection frameworks that is based on the $\chi^{2}$ hypothesis test. However, from a real-world deployment perspective, the sensor data measurements observed in the field might have slight inaccuracies as they are transmitted across the utility ICS. These inaccuracies could arise due to sensor calibration errors and numerical precision limitations especially on inexpensive sensor hardware which are commonly found in large scale ICS. In such cases, expecting to acquire and store a copy of the sensor data measurement at different levels of the ICS that is numerically identical to the one that was originally recorded by the sensor is unrealistic. However given the inherent robustness of hypothesis tests towards robustness, utilizing a sensor data measurement that is similar but not necessarily identical to the originally recorded value is sufficient for establishing statistical consistency. As a result, we develop Lemma 1 as a means to formally establish the notion of probabilistic similarity of sensor measurements and its connection to statistical consistency of the zkSTAR framework.
@@
Probabilistically similar sensor measurement realizations $\tilde{y}_{t},y_{t}$ are necessary and sufficient conditions for observing statistically coherent outcomes $\tilde{\rho}_{t},\rho_{t}$.
@@
Proof provided in Appendix -D
Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes. Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might be able to succeed only when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements. Practically, such attacks end up being inconsequential [22] since they are unable to effect any material damage to the ICS assets in the first place. Therefore, Lemma 1 serves as the basis for Theorem IV.2 which derives an integral link between statistical and temporally consistency on one hand and the key invariance and verification status on the other.
Theorem IV.2 (zk-SNARK Statistical Consistency Conditions).
Given consecutive time step pairs $t,t+1$ with $(vk_{t},pk_{t})$ and $(vk_{t+1},pk_{t+1})$ and a temporally consistent system, statistical consistency can be guaranteed if and only if $\Phi_{t+1}=1$
@@
Proof provided in Appendix -E
@@
Theorem IV.1 shows that alarm values obtained from zkSTAR can be deemed statistically consistent if they share the same proving and verification keys, resulting in a successful verification while adhering to Theorem IV.1. As a result, a regulatory agency can reject the veracity of reported alarms from utility stakeholder in spite of successful verification if either of the conditions laid down in Theorem IV.2 are not met. Additionally, Theorem IV.2 shows that unsuccessful verification of stakeholder proofs implies a statistical inconsistency in alarm detection despite fulfilling temporal consistency claims. Lastly, Theorem IV.2 also establishes that temporal inconsistencies necessarily indicate a statistically inconsistent system regardless of the alarm and verification status.
V Computational Architecture of zkSTAR
@@
The algorithmic framework of zkSTAR is based on a temporal state space modeling paradigm that is driven by a non-linear LSTM-based Kalman filter model. There are three fundamental challenges that arise when attempting to detect data-driven attacks in ICS in zero knowledge. First, asserting temporal consistency between successive state space estimations is challenging due to the highly non-linear and dynamic nature of utility based ICS subsystems. Second, ensuring the statistical consistency of the detection mechanism without revealing parameters of the statistical test is difficult. Lastly, the nature of high-fidelity streaming sensor datasets possess a fine temporal granularity that leads to an additional layer of complexity when characterizing attacks.
As a result, the zkSTAR framework is fundamentally designed to ensure temporal consistency (TC) and statistical consistency (SC) in the context of high-fidelity streaming sensor data. In order to address the challenge of high fidelity sensor information the zkSTAR framework adopts a temporal decomposition scheme that is specifically designed to cater to TC and SC aspects. Therefore, in this section, we provide an in-depth overview of the temporal decomposition scheme before proceeding with a technical discussion and algorithmic review of TC and SC components of zkSTAR.
V-A Temporal Decomposition Scheme
@@
We consider a real-time attack detection horizon that is discretized into a sequence of detection windows denoted by set $\mathcal{W}_{i},\forall i\in\{1,2,\ldots\}$. Each detection window $\mathcal{W}_{i}$ is a set of $W$ time intervals corresponding to utility ICS state space model. Each time interval is represented by $w_{ij},\forall j\in\{1,W\}$ and is primarily geared towards enforcing temporal consistency across multiple prediction time steps. Every TC interval is denoted by $w_{ij},i\in\{1,2,\ldots\},j\in\{1,W\}$ and consists of $D$ consecutive timestamps such that $w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\}$, where $t_{w_{ij}}$ is the starting timestamp of detection window $w_{ij}$. The relationship between each detection window $\mathcal{W}_{i}$ and its constituent TC intervals $w_{ij}$ can be compactly denoted using the following equations.
@@
Without loss of generality, let us consider a detection window $\mathcal{W}$ and its constituent TC intervals $w_{j}\in\mathcal{W},\ j\in\{1,W\}$. We know from Section III that for any time step $t_{w}+k\in w$, $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$ where $r_{t_{w}+k},S_{t_{w}+k}$ are the residual and the covariance matrix predicted by the pre-trained EKF model for time step $t_{w}+k$. Since $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$, the residuals and covariance matrices can be aggregated across all TC intervals for detection window to yield $r_{W}$ and $S_{W}$ as represented in Equation (11).
@@
The temporal decomposition scheme paves the way to leverage zkSNARKs that can help enforce integrity of state space predictions based on a pre-committed EKF model. However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems. Therefore, in order to achieve computational efficiency, while retaining the practicality of zkSNARKs, we adopt a kernel based strategy that decouples compute tasks pertaining to an EKF based state space attack detection framework. Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to temporal consistency of EKF state space predictions as well as statistically consistent outcomes of the detection model that operates on batched EKF residual outputs.
We also note that state of the art zkSNARK frameworks currently do not support linear algebra operations, such as linear solvers, SVD decompositions while lacking the capability of evaluating derivatives of complex sub-modules. This lack of support largely emanates from the significant computational cost of conducting these operations in full zero knowledge making them impractical. In order to circumvent these constraints, we adopt a compute-offline-validate-online approach wherein computationally expensive steps such as derivative evaluations, linear system solutions and SVD decompositions are carried out off-circuit but are evaluated for consistency with the state space model within the zkSNARK circuit.
@@
V-B zkSNARK based Temporal Kernels
For ensuring temporally consistent outcomes, our kernel design strategy involves the breaking down distinct components of the EKF model inference. Specifically, we decompose the EKF inference over a single TC interval into distinct phases that pertain to validating the consistency of the Kalman gain matrix with the state covariance and process and measurement Jacobians; inference of the pre-committed EKF model on one timestamp of the TC interval; and accumulating the residual and its covariance sum across the entire TC interval.
Kalman Reconstruction Check (KRC) Kernel: The EKF model presented in Section III relies on successive linear solves of Equation (28) in order to determine the Kalman gain $K_{t}$. However, this poses a major challenge since implementing linear system solvers within a zkSNARK is computationally very expensive. Therefore, we develop a computational kernel tailored for the zkSNARK paradigm that demonstrates that offline computed matrices Kalman gain $K$, LSTM jacobian $H$, residual covariance $S$ and state covariance $P$ are consistent with respect to the state space EKF model. The result is the KRC kernel depicted in Algorithm 1 which yields a boolean output $\kappa_{t}$ depending on whether the error in Equation (28) is below or above threshold $\epsilon_{KC}$
@@
1:function KRC-Kernel($K$, $S^{-1}$, $H$, $P$, $\epsilon$)
@@
5: compute inverse innovation $S_{t}^{-1}$ using Equation (27).
@@
Algorithm 2 represents the computational kernel for EKF within the zkSTAR framework denoted by EKF-Kernel. For a specific timestamp $t$, EKF-Kernel computes the priori state and covariance estimates $\hat{x}_{t|t-1},P_{t|t-1}$ as well as the resulting residuals $r_{t}$ using Equations (24), (26) and (25) respectively. Using Equation (27), the kernel estimates the inverse of the innovation matrix or the residual covariance $S_{t}^{-1}$. Next, the KRC-Kernel is invoked to validate the consistency of the Kalman gain matrix $K_{t}$ with respect to the Jacobians $G_{t}$, $H_{t}$ as well as the posteriori covariance $P_{t|t}$. Subsequently, the posteriori state and covariance estimates are updated according to Equations (29) and (30). The kernel returns the current residual $r_{t}$, the residual covariance $S_{t}$ and the Kalman gain reconstruction status $\kappa_{t}$.
Temporal Consistency (TC) Kernel: While the EKF kernel can enforce state space dynamics for each timestamp $t$, it cannot independently ensure the integrity of state space estimates across multiple time stamps. In order to do so, we need a mechanism to assert the integrity between distinct timestamps. Therefore, we develop a temporal consistency (TC) kernel that is geared towards ensuring the integrity across dictinct timestamps. Using the temporal decomposition scheme highlighted in Section V-A we focus on a single TC interval denoted by $w$ to aid our discussion. We denote the inputs to the TC kernel in terms of three sets: $\mathcal{X}^{TC}_{w-1},\mathcal{R}^{TC}_{w-1}$ and $\mathcal{K}^{TC}_{w-1}$ which represent state, residual and KRC status estimates from the previous TC interval w1w-1.
@@
Each set in turn consists of the raw values that are appended with a unique randomly chosen none value to provide a good degree of entropy to the resulting hash strings. As a result, $\mathcal{X}^{TC}_{w-1}$ consists of $(x_{w-1|w-1}||N^{x}_{w-1}),(P_{w-1|w-1}||N^{P}_{w-1})$ where $x_{w-1|w-1}$, $P_{w-1|w-1}$, $N^{x}_{w-1}$ and $N^{P}_{w-1}$ denote the posteriori state and covariance estimate and their corresponding nonce values from the previous TC interval. Similarly, $\mathcal{R}^{TC}{w-1}$ comprises the residual and covariance pairs $(r{w-1}||N^{r}_{w-1})$ and $(S_{w-1}||N^{S}_{w-1})$, where $r_{w-1}$ and $S_{w-1}$ represent the innovation and its associated covariance from the previous window, and $N^{r}_{w-1}$ and $N^{S}_{w-1}$ are the corresponding nonce values ensuring stochastic uniqueness of their hash representations. Finally, $\mathcal{K}^{TC}{w-1}$ encapsulates $(\kappa_{w-1}||N^{\kappa}_{w-1})$, where $\kappa{w-1}$ denotes the binary KRC indicator and $N^{\kappa}_{w-1}$ is its appended nonce value used to prevent deterministic linkage across temporal proofs.
@@
Additionally, the TC kernel also utilizes the set $\mathcal{Z}^{TC}_{w}$ which encapsulates the offline computed inputs and sensor data denoted by $\mathcal{Z}^{TC}_{t}$ meant for the EKF Kernel invocations at each timestep of the TC interval. The set $\mathcal{Z}^{TC}_{t}$ contains the nonce-appended sensor measurements and model parameters $(y_{t}||N^{y}_{t})$, $(G_{t}||N^{G}_{t})$, $(H_{t}||N^{H}_{t})$, and $(K_{t}||N^{K}_{t})$ corresponding to the sensor data, the Jacobians, and Kalman gain, respectively. Each element in $\mathcal{Z}^{TC}{t}$ is locally randomized through its associated nonce to ensure that the resulting hashed representations remain unlinkable across successive EKF invocations. Collectively, the windowed set $\mathcal{Z}^{TC}_{w}$ aggregates these per-timestep inputs over the duration of the TC interval $w$, thereby serving as the complete batch of EKF inputs required for evaluating the zkSTAR TC kernel and generating the proof of temporal consistency for that interval. Consequently, Equation (18) is used to represent the TC kernel as a mapping denoted by $\mathcal{M}^{TC}$.
@@
8: accumulate residual sum $r_{w}\leftarrow r_{w}+r_{t}$
9: accumulate residual covariance sum $S_{w}\leftarrow S_{w}+S_{t}$
10: accumulate KRC status $\kappa_{w}\leftarrow\kappa_{w}\circ\kappa_{t}$
11:end for
12:#Append respective nonce strings from $\mathcal{N}^{TC}_{w}$
@@
16:return $\mathcal{X}^{TC}_{w},\mathcal{R}^{TC}_{w},\mathcal{K}^{TC}_{w}$
17:end function
Algorithmically, the kernel is represented using the TC-Kernel function presented in in Algorithm 3. The kernel consumes the outputs of the previous TC interval w1w-1 and iterates through the timestamps $\{t_{w}\ldots t_{w}+D\}$ belonging to the current TC interval $w$. For each timestamp, the kernel strips the nonce vectors to obtain the underlying values which are fed to EKF-Kernel. The residual, covariance and KRC status outputs of the EKF kernel are accumulated across all the timestamps of the current TC interval $w$ through summation and multiplication operations respectively. At each iteration of the kernel, Finally, the kernel returns as the overall outputs of the TC kernel itself.
@@
V-C zkSNARK based Statistical Kernels
Applying the $\chi^{2}$ test on the accumulated residuals and covariance requires us to carry out SVD decomposition to evaluate the test statistic and the alarm criteria. Conducting the SVD decomposition, evaluating the corresponding PC scores, evaluating the test statistic and the subsequent alarm value must be conducted completely in zero knowledge. Therefore, we design a sequence of statistical zkSNARK kernels to help achieve efficiency in zkSTAR.

### v2

@@ 4.2 Temporal Consistency
The state space system is considered to be temporally consistent if for each consecutive time step pairs $t,t+1$, the following holds.
@@
As a consequence of Definition 7 we can make the temporal consistency claim presented in Theorem 4.1.
Theorem 4.1 (zk-STAR TC Conditions).
@@
Proof provided in Appendix B.1
Theorem 4.1 shows that the integrity of temporal posteriori predictions from an EKF driven state space system can be established using a zkSNARK proof artifact if and only if the verification keys remain invariant; the hash commitments of the inputs to the EKF model at each time step is equal to that of the previous EKF output; and the proof artifact itself can be verified using the verification keys. In other words, Theorem 4.1 demonstrates that if temporally inconsistent values are used for carrying out the EKF predictions, then either the verification fails for at least one time step or the keys used for verifying the proof artifact will differ. Therefore, to establish temporal consistency from a regulatory perspective, it is necessary and sufficient to check for the invariance of keys, alignment of hash commitments and the verification status across successive time steps.
4.3 Statistical Consistency
We consider the local attack detection framework characterized by the $\chi^{2}$ hypothesis test presented in Section 3.2. The $\chi^{2}$ test relies on the residual and state space estimates provided by $\mathcal{M}$ that is parametrized by $\theta$. As a result of Equation (8), the local attack detection framework is expected to yield an alarm value denoted by $\rho_{t}=1$, when an attack is taking place while maintaining $\rho_{t}=0$ under normal circumstances. We utilize these fundamental principles to establish the notion of statistical consistency (SC). For the purposes of statistical consistency, we denote the genuine and the reported outcomes of the $\chi^{2}$ test as $\rho^{real}_{t}\in\{0,1\}$ and $\rho^{rep}_{t}\in\{0,1\}$ respectively.
@@ 4.3 Statistical Consistency
Statistical consistency ensures that reported alarm values are consistent with the outcomes of the attack detection frameworks that is based on the $\chi^{2}$ hypothesis test. However, from a real-world deployment perspective, the sensor data measurements observed in the field might have slight inaccuracies as they are transmitted across the utility ICS. These inaccuracies could arise due to sensor calibration errors and numerical precision limitations especially on inexpensive sensor hardware which are commonly found in large scale ICS. In such cases, expecting to acquire and store a copy of the sensor data measurement at different levels of the ICS that is numerically identical to the one that was originally recorded by the sensor is unrealistic. However, given the inherent robustness of hypothesis tests, utilizing a sensor measurements that is similar but not necessarily identical to the originally recorded value is sufficient for establishing statistical consistency. As a result, we develop Lemma 1 as a means to formally establish the notion of probabilistic similarity of sensor measurements and its connection to statistical consistency of the zkSTAR framework.
@@
Probabilistically similar sensor measurement realizations $\tilde{y}_{t},y_{t}$ are necessary and sufficient conditions for observing statistically consistent outcomes $\tilde{\rho}_{t},\rho_{t}$.
@@
Proof provided in Appendix B.2
Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes. Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might in theory go undetected under the zkSTAR framework when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements. Practically, such attacks end up being inconsequential [li2021degradation] since they are unable to effect any material damage to the ICS assets. Therefore, Lemma 1 serves as the basis for Theorem 4.2 which derives an integral link between key invariance, statistical and temporally consistency.
Theorem 4.2 (zk-SNARK SC Conditions).
Given consecutive time step pairs $t,t+1$ with $(vk_{t},pk_{t})$ and $(vk_{t+1},pk_{t+1})$, temporal consistency implies statistical consistency.
@@
Proof provided in Appendix B.3
@@
Theorem 4.2 shows that alarm values obtained from zkSTAR can be deemed statistically consistent if they share the same proving and verification keys and if TC conditions are met according to 4.1. As a result, a regulatory agency can deem reported alarms from utility stakeholder as a violation of SC claims despite successful verification if conditions pertaining to key-invariance or temporal consistency are not met. We formalize the security model with the help of Theorems 4.1 and 4.2 in the subsequent section.
4.4 Security Model and Guarantees
We prove that zkSTAR satisfies knowledge-soundness, guaranteeing that any attempt to forge or reuse stale state estimates for regulatory verification violates the TC-SC consistency relation, thereby ensuring attack-suppression detection. For a given timestep $t$, we denote $\Omega_{t}$ to be the witness for corresponding the state space model $\mathcal{M}$ parametrized by $\theta$. Using Theorems 4.1 and 4.2, the temporal-statistical consistency relation $\mathcal{R}_{t}$ can be defined as in Equation (9) from the perspective of a non-attack.
$\mathcal{R}_{t}\Big[(\Delta_{t}^{in},$ $\Delta_{t}^{out}),\Omega_{t}\Big]=\Big[(\mathcal{M}(\Delta_{t}^{in};\theta)=\Delta_{t}^{out})$ (9)
$\wedge(\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t})\wedge(T_{\chi^{2},t}(\tau_{t})\leq\chi^{2}_{m,\alpha})\Big].$
We consider a utility prover $\mathcal{U}$ undergoing a data-driven cyberattack starting at timestep $t^{attack}>>t^{init}$ where $t^{init}$ represents the initial time step pertaining to the setup phase (Definition 3). For concreteness, we base our analysis on the Halo2 backend for establishing knowledge-soundness. For the analysis of the security model and to establish knowledge-soundness, we assume that the objective of the malicious attacker is to carry out the attack while suppressing attack detection. Moreover, we assume that the compromised utility possesses an extractor $\mathcal{E}^{\mathcal{U}}$ that produces a valid proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{U}}(\Delta_{t}^{in},\Delta_{t}^{out})\neq\Omega_{t}$, with a non-negligible probability. Formally, we define attack suppression using Definition 9.
Definition 9 (Detection Suppression).
The detection framework characterized by $\mathcal{M}(.;\theta)$ is deemed to have been suppressed when the following conditions hold
[…]
$\big((\Delta_{t}^{in},\Delta_{t}^{out}),\,\mathcal{E}^{\mathcal{U}}(\Delta_{t}^{in},\Delta_{t}^{out})\big)\notin\mathcal{R}_{t}$
As a result using the TC and SC claims detailed in Theorems 4.1 and 4.2, we characterize the infeasibility of detection suppression in Theorem 4.3.
Theorem 4.3 (Detection Suppression and Knowledge Soundness).
Given a utility prover $\mathcal{U}$ a timestep $t>t^{attack}$, detection suppression occurs with only negligible probability under the knowledge-soundness property of the underlying proving system. Furthermore, infeasibility of detection suppression is equivalent to violation of either proof verification or of the temporal-statistical consistency relation $\mathcal{R}_{t}$.
Proof.
Proof provided in Appendix B.4. ∎
Theorem 4.3 fundamentally demonstrates that a utility undergoing an attack must either fail proof verification or violate TC-SC relations enforced by zkSTAR. In other words, an under-attack utility cannot fabricate or reuse stale system states without detection, since any such attempt would contradict the knowledge-soundness of the Halo2 backend rendering attack suppression cryptographically infeasible.
5 Computational Architecture of zkSTAR
The zkSTAR architecture is based on the existence of two components comprising the utility as the prover and the regulator as the verifier. As a result, the algorithmic core of zkSTAR focuses on the utility component, encapsulating the utility level EKF model discussed in Section 3 in a zero-knowledge framework. The regulatory component deals with verifying the utility generated proof artifacts and asserting TC, SC and detection suppression claims of utilities.
The framework is designed to generate proofs in an on-demand, lazy fashion using keys generated at the utilities. As a result, utilities commit to the EKF model a priori and use the keys to generate only witnesses in real time containing state space estimates including residuals, covariances and test statistics for each detection window. A regulator can invoke proof generation on-demand at the utility level for specific detection windows as per verification needs.
@@ 5 Computational Architecture of zkSTAR
When attempting to encapsulate the utility level EKF model into a zkSNARK structure, we face three crucial challenges. First, asserting temporal consistency is challenging due to the highly non-linear and dynamic nature of utility based ICS subsystems. Second, ensuring statistically consistent detection outcomes without revealing parameters of the statistical test is difficult. Lastly, high-fidelity streaming sensor datasets possess a fine temporal granularity that leads to an additional layer of computational complexity when characterizing attacks. In order to address these challenges, the zkSTAR framework adopts a temporal decomposition scheme that specifically caters to TC and SC aspects.
5.1 Temporal Decomposition Scheme
@@
We consider a real-time attack detection horizon that is discretized into a sequence of detection windows denoted by set $\mathcal{W}_{i},\forall i\in\{1,2,\ldots\}$. Each detection window $\mathcal{W}_{i}$ is in turn divided into a set of $W$ TC intervals. Each TC interval is represented by $w_{ij},\forall j\in\{1,W\}$ and is primarily geared towards enforcing TC across multiple prediction time steps. Every TC interval consists of $D$ consecutive timestamps such that $w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\}$, where $t_{w_{ij}}$ is the starting timestamp of TC interval $w_{ij}$. The relationship between each detection window $\mathcal{W}_{i}$, its constituent TC intervals $w_{ij}$ and corresponding timesteps can be compactly denoted using Equations (10) and (11).
@@
Without loss of generality, let us consider a detection window $\mathcal{W}$ and its constituent TC intervals $w_{j}\in\mathcal{W},\ j\in\{1,W\}$. We know from Section 3 that for any time step $t_{w}+k\in w$, $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$ where $r_{t_{w}+k},S_{t_{w}+k}$ are the residual and the covariance matrix predicted by the pre-trained EKF model for time step $t_{w}+k$. Since $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$, the residuals and covariance matrices can be aggregated across all TC intervals in a detection window to yield $r_{W}$ and $S_{W}$ denoted in Equation (12).
@@
The temporal decomposition scheme paves the way for zkSNARKs to help enforce integrity of state space predictions based on a pre-committed EKF model. However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems. Therefore, in order to balance practicality and computational efficiency, we adopt a kernel based strategy that decomposes various compute tasks pertaining to an EKF model. Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to TC and SC attributes of EKF predictions and detection outcomes capable of operating on batched EKF residual outputs. We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [ezkl2024]. To address these limitations, we employ a compute-offline–validate-online strategy, performing such expensive operations off-circuit while verifying their consistency with the state-space model within the zkSNARK circuit.
@@
5.2 zkSNARK based Temporal Kernels
For ensuring temporally consistent outcomes, our kernel design strategy involves breaking down distinct components of the EKF model inference. Specifically, we decompose the EKF inference over a single TC interval into distinct phases that pertain to validating the consistency of the Kalman gain matrix with the state covariance and process and measurement Jacobians; inference of the pre-committed EKF model on one timestamp of the TC interval; and accumulating the residual and its covariance sum across the entire TC interval.
Kalman Reconstruction Check (KRC) Kernel: The EKF model presented in Section 3 relies on successive linear solves of Equation (5) in order to determine the Kalman gain $K_{t}$. However, this poses a major challenge since implementing linear system solvers within a zkSNARK is computationally very expensive. Therefore, we develop a computational kernel tailored for the zkSNARK paradigm that demonstrates that offline computed matrices Kalman gain $K$, LSTM Jacobian $H$, residual covariance $S$ and state covariance $P$ are consistent with respect to the state space EKF model. The result is the KRC kernel depicted in Algorithm 1 which yields a boolean output $\kappa_{t}$ depending on whether the error in Equation (5) is below or above $\epsilon_{KC}$
@@
1:function KRC-Kernel($K$, $S$, $H$, $P$, $\epsilon$)
@@
5: compute innovation $S_{t}$ using Equation (4).
@@
Algorithm 2 represents the computational kernel for EKF within the zkSTAR framework denoted by EKF-Kernel. For a timestamp $t$, EKF-Kernel computes the priori state and covariance estimates $\hat{x}_{t|t-1},P_{t|t-1}$ as well as the resulting residuals $r_{t}$ using Equations (1), (3) and (2) respectively. Using Equation (4), the kernel estimates the innovation or the residual covariance matrix $S_{t}$. Next, the KRC-Kernel is invoked to validate the consistency of the Kalman gain matrix $K_{t}$ with respect to the Jacobian $H_{t}$, the residual covariance $S_{t}$, as well as the posteriori covariance $P_{t|t}$. Subsequently, the posteriori state and covariance estimates are updated according to Equations (6) and (7). The kernel returns the current residual $r_{t}$, the residual covariance $S_{t}$ and the Kalman reconstruction status $\kappa_{t}$.
Temporal Consistency (TC) Kernel: While the EKF kernel can enforce state space dynamics for each timestamp $t$, it cannot independently ensure the integrity of state space estimates across multiple time stamps. In order to do so, we need a mechanism to assert the integrity between distinct timestamps. Therefore, we develop a temporal consistency (TC) kernel that is geared towards ensuring the integrity across distinct timestamps. Using the temporal decomposition scheme highlighted in Section 5.1 we focus on a single TC interval denoted by $w$ to aid our discussion. We denote the inputs to the TC kernel in terms of three sets: $\mathcal{X}^{TC}_{w-1},\mathcal{R}^{TC}_{w-1}$ and $\mathcal{K}^{TC}_{w-1}$ which represent state, residual and KRC status estimates from the previous TC interval w1w-1.
@@
Each set in turn consists of the raw values that are appended with a unique randomly chosen nonce value to provide a robust degree of entropy to the resulting hash strings. As a result, $\mathcal{X}^{TC}_{w-1}$ consists of $(x_{w-1|w-1}||N^{x}_{w-1}),(P_{w-1|w-1}||N^{P}_{w-1})$ where $x_{w-1|w-1}$, $P_{w-1|w-1}$, $N^{x}_{w-1}$ and $N^{P}_{w-1}$ denote the posteriori state and covariance estimate and their corresponding nonce values from the previous TC interval. Similarly, $\mathcal{R}^{TC}_{w-1}$ comprises the residual and covariance pairs $(r{w-1}||N^{r}_{w-1})$ and $(S_{w-1}||N^{S}_{w-1})$, where $r_{w-1}$ and $S_{w-1}$ represent the innovation and its associated covariance from the previous window, and $N^{r}_{w-1}$ and $N^{S}_{w-1}$ are the corresponding nonce values ensuring stochastic uniqueness of their hash representations. Finally, $\mathcal{K}^{TC}_{w-1}$ encapsulates $(\kappa_{w-1}||N^{\kappa}_{w-1})$, where $\kappa_{w-1}$ denotes the binary KRC indicator and $N^{\kappa}_{w-1}$ is its appended nonce value used to prevent deterministic linkage across temporal proofs.
@@
Additionally, the TC kernel also utilizes the set $\mathcal{Z}^{TC}_{w}$ which encapsulates the offline computed inputs and sensor data denoted by $\mathcal{Z}^{TC}_{t}$ meant for the EKF Kernel invocations at each timestep of the TC interval. The set $\mathcal{Z}^{TC}_{t}$ contains the nonce-appended sensor measurements and model parameters $(y_{t}||N^{y}_{t})$, $(G_{t}||N^{G}_{t})$, $(H_{t}||N^{H}_{t})$, and $(K_{t}||N^{K}_{t})$ corresponding to the sensor data, the Jacobians, and Kalman gain, respectively. Each element in $\mathcal{Z}^{TC}_{t}$ is locally randomized through its associated nonce to ensure that the resulting hashed representations remain unlinkable across successive EKF invocations. Collectively, the windowed set $\mathcal{Z}^{TC}_{w}$ aggregates these per-timestep inputs over the duration of the TC interval $w$, thereby serving as the complete batch of EKF inputs required for evaluating the zkSTAR TC kernel and generating the proof of temporal consistency for that interval. Consequently, Equation (19) is used to represent the TC kernel as a mapping denoted by $\mathcal{M}^{TC}$.
@@
[…]
10: accumulate residual sum $r_{w}\leftarrow r_{w}+r_{t}$
11: accumulate residual covariance sum $S_{w}\leftarrow S_{w}+S_{t}$
12: accumulate KRC status $\kappa_{w}\leftarrow\kappa_{w}\circ\kappa_{t}$
13:end for
14:#Append nonce strings from $\mathcal{N}^{TC}_{w}$
@@
18:return $\mathcal{X}^{TC}_{w},\mathcal{R}^{TC}_{w},\mathcal{K}^{TC}_{w}$
19:end function
Algorithmically, the kernel is represented using the TC-Kernel function presented in Algorithm 3. The kernel consumes the outputs of the previous TC interval w1w-1 and iterates through the timestamps $\{t_{w}\ldots t_{w}+D\}$ belonging to the current TC interval $w$. For each timestamp, the kernel strips the nonce vectors to obtain the underlying values which are fed to EKF-Kernel. The residual, covariance and KRC status outputs of the EKF kernel are accumulated across all the timestamps of the current TC interval $w$ through summation and multiplication operations respectively. At each iteration, the kernel returns the overall outputs pertaining to the state, residual and reconstruction status check for the corresponding TC interval.
@@
5.3 zkSNARK based Statistical Kernels
Applying the $\chi^{2}$ test on the accumulated residuals and covariance requires us to carry out SVD decomposition to evaluate the test statistic and the alarm criteria in complete zero-knowledge. Therefore, we design a sequence of statistical zkSNARK kernels to help achieve efficiency in zkSTAR.